40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13749     Entry time: Thu Apr 12 18:12:49 2018     In reply to: 13746     Reply to this: 13751
Author: gautam 
Type: Update 
Category: ALS 
Subject: NPRO channels hijacked 

Summary:

  1. Today, the measured IR ALS noise for the X arm was dramatically improved. The main change was that I improved the alignment of the PSL pickoff beam into its fiber coupler.
  2. The noise level was non-stationary, leading me to suspect power modulation of the RF beat amplitude.
  3. I am now measuring the stability of the power in the two polarizations coming from EX table to the PSL table, using the PSL diagnostic connector channels.
  4. The EX beam is S-polarized when it is coupled into the fiber. The PSL beam is P-polarized. However, it looks like I have coupled light along orthogonal axes into the fiber, such that when the EX light gets to the PSL table, most of it is in the P-polarization, as judged by my PER measurement setup (i.e. the alignment keys at the PSL table and at the EX table are orthogonal). So it still seems like there is something to be gained by trying to improve the PER a bit more.

Details:

Today, I decided to check the power coupled into the PSL fiber for the BeatMouth. Surprisingly, it was only 200uW, while I had ~3.15mW going into it in January. Presumably some alignment drifting happened. So I re-aligned the beam into the fiber using the steering mirror immediately before the fiber coupler. I managed to get ~2.9mW in without much effort, and I figured this is sufficient for a first pass, so I didn't try too much more. I then tried making an ALS beat spectrum measurement (arm locked to IMC length using POX, green following the arm using end PDH servo). Surprisingly, the noise performannce was almost as good as the reference! See Attachment #1, in which the red curve is an IR beat (while all others are green beats). The Y arm green beat performance isn't stellar, but one problem at a time. Moreover, the kind of coherence structure between the arm error signal and the ALS beat signal that I reported here was totally absent today.

Upon further investigation, I found that the noise level was actually breathing quite significantly on timescales of minutes. While I was able to successfully keep the TEM00 mode of the PSL beam resonant inside the arm cavity by using the ALS beat frequency as an error signal and MC2 as a frequency actuator, the DC stability was very poor and TRX was wandering around by 50%. So my new hypothesis is that the excess ALS noise is because of one or more of

  • Beam jitter at coupling point into fiber.
  • Polarization drift of the IR beams.

While I did some work in trying to align the PSL IR pickoff into the fiber along the fast (P-pol) axis, I haven't done anything for the X end pickoff beam. So perhaps the fluctuations in the EX IR power is causing beatnote amplitude fluctuations. In the delay line + phase tracker frequency discriminator, I think RF beatnote amplitude fluctuations can couple into phase noise linearly. For such an apparently important noise source, I can't remeber ever including it in any of the ALS noise budgets.

Before Ph237 today, I decided to use my polarization monitoring setup and check the "RIN" of power in the two polarizations coming out of the fiber on the PSL table. For this purpose, I decided to hijack the Acromag channels used for the PSL diagnostics connector Attachment #2 shows that there is fluctuations at the level of ~10% in the p-polarization. This is the "desired" polarization in that I aligned the PSL beam into the fiber to maximize the power in this polarization. So assuming the power fluctuations in the PSL beam are negligible, this translates to sqrt(10) ~3% fluctuation in the RF beat amplitude. This is at best a conservative estimate, as in reality, there is probably more AM because of the non PM fibers inside the beatmouth.

All of this still doesn't explain the coherence between the measured ALS noise and the arm error signal at 100s of Hz (which presumably can only happen via frequency noise in the PSL).

Another "mystery" - yesterday, while I was working on recovering the Y arm green beat signal on the PSL table, I eventually got a beat signal that was ~20mVpp into 50ohms, which is approximately the same as what I measured when the Y arm ALS performance was "nominal", more than a year ago. But while viewing the Y arm beats (green and IR) simultaneously on an o'scope, I wasn't able to keep both signals synchronised while triggering on one (even though the IR beat frequency was half the green beat frequency). This means there is a huge amount of relative phase noise between the green and IR beats. What (if anything) does this mean? The differential noise between these two beats should be (i) phase noise at the fiber coupler/ inside the fiber itself, and (ii) scatter noise in the green light transmitted through the cavity. Is it "expected" that the relative phase noise between these two signals is so large that we can't view both of them on a common trigger signal on an o'scope? surpriseAlso - the green mode-matching into the Y arm is abysmal.

Anyways - I'm going to try and tweak the PER and mode-matching into the X end fiber a little and monitor the polarization stability (nothing too invasive for now, eventually, I want to install the new fiber couplers I acquired but for now I'll only change alignment into and rotation of the fiber coupler on the EX table). It would also be interesting to compare my "optimized" PSL drift to the unoptimized EX power drift. So the PSL diagnostic channels will not show any actual PSL diagnostic information until I plug it back in. But I suspect that the EPICS record names and physical channel wiring are wrong anyways - I hooked up my two photodiode signals into what I would believe is the "Diode 1 Power" and "Laser crystal temperature" monitors (as per the schematic), but the signals actually show up for me in "Diode 2 Power" (p-pol) and "Didoe 1 Temperature" (s-pol).

Annoyingly, there is no wiring diagram - on my todo list i guess...

@Steve - could you please take a photo of the EX table and update the wiki? I think the photo we have is a bit dated, the fiber coupler and transmon PDs aren't in it...

Attachment 1: IR_ALS_20180412.pdf  73 kB  Uploaded Thu Apr 12 19:58:15 2018  | Hide | Hide all
IR_ALS_20180412.pdf
Attachment 2: BeatMouthDrift.png  40 kB  Uploaded Thu Apr 12 19:59:11 2018  | Hide | Hide all
BeatMouthDrift.png
Attachment 3: ETMX_20180416.jpg  4.499 MB  Uploaded Mon Apr 16 11:01:40 2018  | Hide | Hide all
ETMX_20180416.jpg
ELOG V3.1.3-